### v1

# Interpretability Illusions with Sparse Autoencoders: Evaluating Robustness of Concept Representations

Abstract
Sparse autoencoders (SAEs) are commonly used to interpret the internal activations of large language models (LLMs) by mapping them to human-interpretable concept representations. While existing evaluations of SAEs focus on metrics such as the reconstruction-sparsity tradeoff, human (auto-)interpretability, and feature disentanglement, they overlook a critical aspect: the robustness of concept representations to input perturbations. We argue that robustness must be a fundamental consideration for concept representations, reflecting the fidelity of concept labeling. To this end, we formulate robustness quantification as input-space optimization problems and develop a comprehensive evaluation framework featuring realistic scenarios in which adversarial perturbations are crafted to manipulate SAE representations. Empirically, we find that tiny adversarial input perturbations can effectively manipulate concept-based interpretations in most scenarios without notably affecting the outputs of the base LLMs themselves. Overall, our results suggest that SAE concept representations are fragile and may be ill-suited for applications in model monitoring and oversight. The code for our experiments is available at https://github.com/AI4LIFE-GROUP/sae_robustness.
1.  Introduction
As large language models (LLMs) become widely used across diverse applications, the need to monitor and summarize their internal representations is critical for both interpretability and reliability. Sparse autoencoders (SAEs) (Cunningham et al., 2023) have shown promise as an unsupervised approach to map LLM embeddings to sparse interpretable concept embeddings via dictionary learning, where each neuron’s activation can be associated with specific, human-understandable concepts. Besides the reconstruction-sparsity Pareto frontier (Gao et al., 2024) and the human-understandability of the learned SAE latents (Paulo et al., 2024), a growing number of recent works have considered SAE’s feature disentanglement and concept detection capabilities (Karvonen et al., 2024a, 2025) as important components in SAE evaluation.
However, while existing works show promise with the usage of SAEs under co-operative contexts, where both the explanation provider and the user share similar incentives; their applicability to adversarial contexts remains underexplored. We borrow the nomenclature of "co-operative" and "adversarial" contexts from (Bordt et al., 2022), who define an adversarial context as one where the model explainer has conflicting goals with the consumer of the explanation. For example, an adversarial user may craft prompts that manipulate SAE activations to bypass refusal mechanisms or produce benign-looking interpretations, thereby evading safety systems built on top of the model’s SAE representations. More broadly, if SAE-derived latent spaces are vulnerable to minimal input perturbations, adversaries could exploit this to conceal harmful, deceptive, or biased model outputs from downstream users. Conversely, if minor variations in inputs lead semantically unrelated prompts to yield similar SAE representations, it is challenging to assign precise, human-interpretable labels to individual concepts and to reason systematically about the effects of representation editing, leading to questions about the fragility of concept labeling. These vulnerabilities extend beyond what prior evaluations of monosemanticity (Bricken et al., 2023; Minegishi et al., 2025), which assume static, coherent, and unperturbed inputs, could uncover, raising deeper concerns about the robustness and reliability of SAE interpretations in trustworthy AI contexts.
In this work, we propose incorporating robustness as a core component of the SAE evaluation agenda, and concept-based explanations in general. We begin by defining SAE robustness under input-level perturbations, followed by the specification of adversarial settings for a comprehensive robustness evaluation. Each setting is characterized by two key objectives and a perturbation mode, which determines how the input-level attack is applied. First, the semantic goal: the perturbation could be either untargeted, aiming to alter the SAE activations from the original, or targeted, aiming to match the perturbed activations to that of another semantically unrelated input. Second, the activation goal: either at the population level, where the objective is to manipulate all SAE neurons simultaneously to alter the overall activation pattern; or at the individual level, where the goal is to activate a specific feature independently. Additionally, we consider two basic attack modes in this work: suffix-based attacks, where new tokens are appended to the input, and replacement-based attacks, where original tokens are allowed to be substituted with adversarial ones. Putting these together, we define eight different scenarios to evaluate SAE concept robustness against adversarial perturbations.
To search for adversarial inputs that manipulate SAE interpretations, we adapt Greedy Coordinate Gradient (GCG) (Zou et al., 2023), a widely used input-space search algorithm originally developed to generate adversarial prompts that induce harmful outputs in LLMs, to our SAE settings. Our findings reveal that current SAEs are not sufficiently robust to serve as reliable tools for interpreting and monitoring LLM behaviors, nor as trustworthy components in downstream applications that depend on stable, concept-level representations. We further believe that this vulnerability to adversarial input manipulation could generalize to other SAE-inspired approaches such as transcoders (Dunefsky et al., 2024) and crosscoders (Lindsey et al., 2024), highlighting the need for future research on developing more robust and reliable concept-extraction methods in the field of mechanistic interpretability.
Our main contributions can be summarized as follows:
  • We identify robustness as a critical yet underexplored dimension in evaluating SAEs, expanding the current evaluation agenda by introducing input-level perturbations.
  • We propose a comprehensive evaluation framework that defines SAE robustness along semantic goals, activation goals, and perturbation modes, resulting in eight distinct evaluation scenarios.
  • We conduct extensive experiments by designing adversarial input-level attacks, showing that SAE interpretations are consistently vulnerable to input perturbations across different datasets and model configurations.
2.  Related Work
SAE as an Interpretability Tool
Since SAE was first proposed by Cunningham et al. (2023) as an effective approach for mechanistic interpretability (Bereska and Gavves, 2024; Sharkey et al., 2025), extensive works have focused on improving its architectural design (Rajamanoharan et al., 2024a; Mudide et al., 2024), activation functions (Gao et al., 2024; Rajamanoharan et al., 2024b; Bussmann et al., 2024), and loss functions (Karvonen et al., 2024b; Marks et al., 2024a). SAEs have been applied to study LLM internal dynamics (Kissane et al., 2024; Ziyin et al., 2024; O’Neill et al., 2024; Balagansky et al., 2024; Lawson et al., 2024), control model behaviors (Marks et al., 2024b; Chalnev et al., 2024), as well as facilitate various downstream applications (Magalhães et al., 2024; Lei et al., 2024).
Evaluation of SAEs
Beyond the reconstruction–sparsity tradeoff (Gao et al., 2024), which has largely shaped the design of SAE training objectives, and the alignment of learned latents with human knowledge (i.e., human-understandability) (Cunningham et al., 2023; Paulo et al., 2024), recent works have begun to assess SAE performance from a more interpretation-centric perspective (Makelov et al., 2024; Karvonen et al., 2025; Bhalla et al., 2024). These efforts include evaluating whether prespecified, meaningful concepts can be captured by individual latents (Gurnee et al., 2023; Chanin et al., 2024) and whether independent semantic features are properly disentangled in the latent space (Huang et al., 2024; Karvonen et al., 2024a). Our work complements and extends these static evaluations of concept detection and feature disentanglement (Karvonen et al., 2025) by introducing adversarial perturbations at the input level to assess the robustness of SAE-derived interpretations.
Adversarial Attacks and Prompt Optimization
LLMs are known to be vulnerable to adversarial attacks in the input space (Chen et al., 2022; Zou et al., 2023; Kumar et al., 2023; Zeng et al., 2024; Das et al., 2025), where small perturbations to prompts can lead to degraded cognitive performance or harmful model generations. Greedy Coordinate Gradient (GCG) (Zou et al., 2023) is a universal prompt optimization paradigm that searches for promising tokens to minimize a specified language model loss. In this work, we generalize GCG to the SAE setting to construct effective adversarial inputs that render SAE interpretations unreliable.
3.  Evaluating the Robustness of SAE Interpretations
In this section, we introduce a formal framework for evaluating SAE robustness. We begin by formulating robustness as input-space optimization problems, then present an evaluation framework based on structured adversarial scenarios, and finally propose a generalized input-level attack algorithm for solving the optimization objectives.
3.1 Preliminaries
Sparse autoencoders (SAEs) are linear layers typically trained on the residual stream of LLMs, with distinct weights for each layer. Formally, the target LLM $f_{LLM}:\mathcal{X}\rightarrow\mathcal{H}$ first maps an input sequence $x$ to a hidden state $h$, and then the SAE $f_{SAE}:\mathcal{H}\rightarrow\mathcal{Z}$ projects it to the sparse latent space. The SAE encoding and decoding processes are given by:
$z=\phi(W_{enc}h+b_{enc})$ (1)
$\hat{h}=W_{dec}z+b_{dec}$ (2)
During encoding, $\phi$ is a sparsity-encouraging activation function, and popular choices include ReLU (Cunningham et al., 2023) and TopK (Gao et al., 2024). During decoding, $\hat{h}$ can be reconstructed as a sparse linear combination of interpretable concepts in the dictionary with a bias term.
3.2 Proposed Theory of SAE Robustness
Independent from LLMs and SAEs, we assume the existence of a ‘ground-truth’ concept mapping $f_{c}$ from the input space $\mathcal{X}$ to a semantic concept space $\mathcal{C}$, such that an input sequence corresponds to a ground truth concept vector capturing the magnitudes of all semantic features. This mapping provides an external reference for interpretability: under this setup, evaluating SAE interpretability can be ultimately considered as assessing the degree of alignment between the learned sparse latent space $\mathcal{Z}$ and $\mathcal{C}$. Ideally, this mapping should be close to a bijection, suggesting both monosemanticity (i.e. a single latent encodes a single concept) and concept identifiability (i.e. a single concept can be captured by a small number of latents) (Karvonen et al., 2025). We now consider the conditions under which such alignment is violated. This can be formally expressed in two directions as:
$\exists c_{1},c_{1}^{\prime}\in\mathcal{C}:d_{c}(c_{1},c_{1}^{\prime})<\epsilon_{c},d_{z}(z_{1},z_{1}^{\prime})>\delta_{z}$ (3)
$\exists c_{1},c_{2}\in\mathcal{C}:d_{c}(c_{1},c_{2})>\delta_{c},d_{z}(z_{1},z_{2})<\epsilon_{z}$ (4)
where $z_{i}=(f_{SAE}\circ f_{LLM}\circ f_{c}^{-1})(c_{i}),\forall c_{i}\in\mathcal{C}$. The distance metrics $d_{c}$ and $d_{z}$, along with the thresholds $\epsilon_{c},\delta_{c},\epsilon_{z},\delta_{z}$, should be chosen based on the criteria for when two semantic concept vectors or SAE activations are considered highly similar or entirely unrelated. Simply stated, the violations holds when similar semantic concepts correspond to dissimilar SAE concept vectors, or vice versa.
However, since directly searching in the hypothetical concept space $\mathcal{C}$ is impractical, existing works (Gurnee et al., 2023; Huang et al., 2024; Marks et al., 2024b; Karvonen et al., 2024a) often simplify this problem setup by focusing on prespecified concepts, and then applying perturbations in $\mathcal{Z}$ with predetermined directions and step sizes. In this setup, $d_{z}$ is typically measured based on the overlap between two sets of top-k activated SAE latents, while $d_{c}$ is quantified by the accuracies of external probes trained to predict the presence of specific concepts.
As we cannot directly apply perturbations in $\mathcal{C}$, we instead propose to apply input perturbations in $\mathcal{X}$, by making a Lipschitz-ness assumption regarding the ‘ground-truth’ concept map $f_{c}$. Specifically, the distance metric $d_{x}$ is defined as the Levenshtein distance (Levenshtein et al., 1966) between two token sequences (i.e. the minimum number of single-token insertions, deletions, or replacements required to transform $x_{i}$ into $x_{j}$), which locally and proportionally reflects the semantic distance $d_{c}$ in the concept space.
Assumption 1.
$f_{c}$ is bi-Lipschitz, i.e.
$L_{1}\cdot d_{x}(x_{i},x_{j})\leq d_{c}(f_{c}(x_{i}),f_{c}(x_{j}))\leq L_{2}\cdot d_{x}(x_{i},x_{j})$
for some constants $L_{1},L_{2}>0$ and all $x_{i},x_{j}$ $\in\mathcal{X}$.
This assumption is motivated by the observation that small changes in inputs typically induce only slight and gradual shifts in overall semantic meaning, which makes token-level edit distance a practical proxy for semantic variations, enabling small perturbations at the concept level without requiring direct access to the hypothetical concept space $\mathcal{C}$. Compared to prior approaches, input-level perturbations offer three distinct advantages:
  • They support more fine-grained control, allowing perturbations in arbitrary directions and with variable step sizes in the input space (as explained in Section 3.3).
  • They enable concept-level evaluations without relying on hand-crafted latent directions or predefined concept labels.
  • They better reflect realistic threat models, as it is significantly easier for an adversary to manipulate raw inputs than to intervene in latent or activation spaces.
Therefore, the search problem in the concept space can be transformed into an optimization problem in the input space, by directly investigating the mapping $f_{LLM}\circ f_{SAE}:\mathcal{X}\rightarrow\mathcal{Z}$. We define the extent to which this bijection is preserved under adversarial input-level perturbations as the robustness of the SAE. For any given input $x_{1}$, this can be quantified by:
$\max_{x_{1}^{\prime}}d_{z}(z_{1},z_{1}^{\prime})\quad\text{subject to}\quad d_{x}(x_{1},x_{1}^{\prime})\leq\epsilon_{x}$ (5)
$\min_{x_{2}}d_{z}(z_{1},z_{2})\quad\text{subject to}\quad d_{x}(x_{1},x_{2})\geq\delta_{x}$ (6)
These two objectives form the foundation of our evaluation framework.
3.3 Proposed Evaluation Framework
Based on the preceding definition of SAE robustness, we propose a structured evaluation framework that further specifies the optimization problem for empirical analysis. The framework consists of three independent binary dimensions: semantic goal, activation goal, and perturbation mode. Each of the eight resulting combinations defines a unique adversarial scenario corresponding to a well-defined optimization task in the input space.
3.3.1 Semantic Goal
The semantic goal determines the direction of the perturbation:
Untargeted
Given $x_{1}$ and a fixed edit distance $\epsilon_{x}$ between token sequences, the attack aims to find a perturbed $x_{1}^{\prime}$ that maximizes the difference in SAE activation. The perturbation direction is not predefined but is empirically selected to induce the maximal change in the sparse latent space $\mathcal{Z}$. This setting corresponds to objective (5) exactly.
Targeted
Given both $x_{1}$ and an entirely unrelated $x_{2}$ as the target, our goal becomes searching for an $x_{1}^{\prime}$ that remains close to $x_{1}$ in $\mathcal{X}$ while resembling $x_{2}$ in $\mathcal{Z}$:
$\min_{x_{1}^{\prime}}d_{z}(x_{1}^{\prime},x_{2})\quad\text{subject to}\quad d_{x}(x_{1},x_{1}^{\prime})\leq\epsilon_{x}$ (7)
This is equivalent to objective (6), since $d_{x}(x_{1},x_{1}^{\prime})\leq\epsilon_{x}$ implies $d_{x}(x_{1}^{\prime},x_{2})\geq\delta_{x}$. While it may seem that the pair $(x_{1},x_{1}^{\prime})$ could also satisfy the objective (5) simultaneously, we nevertheless define the untargeted setting separately, as the perturbation here follows a fixed direction toward $x_{2}$, making it a more constrained scenario.
These two settings capture distinct adversarial objectives: untargeted perturbations evaluate the general fragility of the sparse latent space $\mathcal{Z}$, while targeted perturbations test whether SAE activations can be deliberately steered toward misleading interpretations.
3.3.2 Activation Goal
The activation goal defines the distance metric $d_{z}$ within the SAE latent space. Given two SAE activation vectors $z_{i},z_{j}\in\mathcal{Z}$, the perturbation goal could vary in granularity. In this work, we consider the following two levels:
Population level
The goal is to manipulate groups of SAE features simultaneously to control the overall sparse representation vector, so the distance metric can be defined by the overlap ratio between two sets of $k$ most activated SAE latents (we call this metric neuron overlap ratio throughout this work):
$d_{z}(z_{i},z_{j})=1-\frac{|\mathcal{I}_{k}(z_{i})\cap\mathcal{I}_{k}(z_{j})|}{k}$ (8)
where $\mathcal{I}_{k}(z)$ denotes the set of indices of the $k$ most activated latents in vector $z$. To align with the notion of activating/deactivating SAE latents, $k$ is always set to the number of non-zero latents activated by the target sequence $x_{2}$.
Individual level
The perturbation may also be directed toward a specific SAE feature by modifying its rank among all latent dimensions. In the case of activation, the objective is to increase the feature’s rank until it has non-zero activation after $\phi$. Conversely, in the case of deactivation, the objective is to reduce the feature’s rank such that the latent becomes zero. We thus define $d_{z}$ in a binary manner:
$d_{z}(z_{i},z_{j})=\mathds{1}\left[\mathds{1}_{z_{i}^{(t)}>0}\neq\mathds{1}_{z_{j}^{(t)}>0}\right]$ (9)
where $t$ is the index of the target SAE feature to be manipulated.
These two levels capture complementary structural vulnerabilities of SAE representations. The population-level objective assesses global interpretation shifts, while the individual-level objective probes the microscopic stability of specific SAE features.
3.3.3 Perturbation Mode
Analogous to adversarial attacks on LLMs, the perturbation mode defines how adversarial edits are applied to the input. In this work, we consider two fundamental types:
Suffix mode
Adversarial tokens are appended to the end of the original input. This preserves the original semantic content while allowing the attacker to introduce new features that can dominate the existing ones by appending a limited-length suffix.
Replacement mode
Tokens in the original input are substituted with adversarial ones. Compared to the suffix mode, this approach directly modifies existing semantic features, and we therefore restrict the perturbation to a single token replacement to ensure that the change at the input level remains minimal. The overall robustness under replacement mode can be measured by the average across all token positions.
These two modes represent different types of adversarial control, and both are highly plausible in practical deployment scenarios. By systematically combining semantic goal, activation goal, and perturbation mode, we have defined a total of eight different settings to evaluate SAE robustness.
3.4 Generalized Input-level Attack for SAE
Inspired by Gradient Coordinate Gradient (GCG) (Zou et al., 2023) used to elicit harmful LLM outputs, we propose a generalized algorithm to find best adversarial input-level perturbations, as part of our evaluation framework for SAE robustness.
To search for promising tokens in the discrete input space $\mathcal{X}$, traditional GCG employs an iterative optimization procedure: at each iteration, it first computes gradients with respect to token embeddings using a designated loss function, which is typically a language modeling loss aimed at aligning outputs with expected behavior; it then samples a batch of adversarial prompts based on the gradients, evaluates them under the same loss function, and finally selects the most effective candidate to proceed to the next iteration.
A primary challenge in directly applying GCG to our SAE setting lies in the non-differentiability of the distance metrics in $\mathcal{Z}$, as defined by equations (8) and (9). Therefore, we compute the gradients with differentiable loss functions defined over the continuous SAE representation space, while candidate solutions are evaluated using the original non-differentiable distance metrics defined over the sparse latent space. We summarize the various loss functions and evaluation metrics used for different semantic and activation goals in Table 1.
Activation Goals
Semantic Goals Population Level Individual Level
GCG Loss Evaluation GCG Loss Evaluation
Untargeted $\frac{\tilde{z}_{1}\cdot\tilde{z}_{1}^{\prime}}{\|\tilde{z}_{1}\|\,\|\tilde{z}_{1}^{\prime}\|}$ $\frac{|\mathcal{I}_{k}(z_{1})\cap\mathcal{I}_{k}(z_{1}^{\prime})|}{k}$ $\pm\log\left(\frac{\exp(z_{1}^{\prime(t)})}{\sum_{j}\exp(z_{1}^{\prime(j)})}\right)$ $\pm\operatorname{rank}(z_{1}^{\prime(t)})$
Targeted $\frac{\tilde{z}_{1}^{\prime}\cdot\tilde{z}_{2}}{\|\tilde{z}_{1}^{\prime}\|\,\|\tilde{z}_{2}\|}$ $1-\frac{|\mathcal{I}_{k}(z_{1}^{\prime})\cap\mathcal{I}_{k}(z_{2})|}{k}$ $\pm\log\left(\frac{\exp(z_{1}^{\prime(t)})}{\sum_{j}\exp(z_{1}^{\prime\prime(j)})}\right)$ $\pm\operatorname{rank}(z_{1}^{\prime(t)})$
Table 1: Customized GCG loss functions and evaluation metrics for different combinations of semantic and activation goals. When evaluating individual SAE features, both activation and deactivation tasks are tested.
Here, $\tilde{z}=W_{enc}h+b_{enc}$ denotes the raw activation vector prior to applying the sparsity-inducing activation function. We use cosine similarity and log-likelihood as loss functions in the continuous representation space, while retaining the original distance measures in $\mathcal{Z}$ as criteria for selecting adversarial candidates. The only exception is at the individual level, where we replace the original binary distance metric with the rank of the specified SAE feature.
The complete pseudocode for our generalized input-level attack is provided in Appendix A.
4.  Experimental Evaluation
In this section, we first describe our experimental setup, including model and dataset choices, followed by results across all eight adversarial scenarios. We then present additional analyses to validate the reliability of our evaluation framework, examine robustness trends across model depths, assess cross-model attack transferability, and conclude with a case study on manipulating highly interpretable SAE latents.
4.1 Experimental Setup
Models
We evaluate SAE robustness on two combinations of open-source LLMs and pretrained SAEs on the residual streams of the models. In this section, we present the results for (1) layer 20 of Llama-3-8B with TopK as the SAE activation function (Gao et al., 2024) and (2) layer 30 of Gemma-2-9B with JumpReLU SAE (Rajamanoharan et al., 2024b). Both SAEs have a width of 131k. Our experiments are not conducted on smaller models or SAEs with fewer latents, as our targeted attacks require a relatively large initial distance $d_{z}(x_{1},x_{2})$ to meaningfully evaluate their effectiveness. In other words, we evaluate the robustness of an SAE only when it can clearly distinguish semantically unrelated inputs. We focus on mid-to-late layers, as they strike a balance between low-level surface patterns in early layers and highly entangled representations in final layers, making them better suited for capturing human-interpretable concepts. Additional results for different model depths are included in Appendix F.
Datasets
The ideal evaluation for targeted tasks should be conducted on two sets of input prompts with minimal semantic overlap. Our first ChatGPT-generated Art & Science dataset consists of 100 pairs of sentences each with approximately 20 to 25 tokens, broadly divided into two categories: $x_{1}$ is related to art and humanity, and $x_{2}$ is about science and technology. We adapt another existing dataset AG News to sample 300 pairs of news titles with different category labels as an additional evaluation dataset. The mean initial overlap ratios are $31.3\%$ and $33.4\%$ for the two datasets, averaged across two models.
Evaluation Configurations
Since the residual streams of LLMs encode both semantic features and next-token prediction information, we append a short instruction prompt to the original sequence (including the suffix), "The previous sentence is about", to the input to better extract LLM’s semantic content from the last hidden state. At individual level, we focus on $10$ individual neurons, which are selected for a activation/deactivation task based on the semantic goal: for an untargeted task, we select the SAE latents with the lowest/highest activation values for $x_{1}$; for a targeted task, we select the latents most/least activated by $x_{2}$ but are currently deactivated/activated by $x_{1}$. The specific choices of hyperparameters are included in Appendix B.
Model Semantic Goal Population Level Individual Level
Neuron Overlap (%) Attack Success Rate (%)
Suffix Replacement Suffix Replacement
Llama 3 8B Untargeted $89.1\pm 2.3$ $84.6\pm 0.9$ $\pm 0.0,22.9\pm 1.5 0.0±0.0$$\pm 0.0,34.3\pm 0.9$
Targeted $74.5\pm 3.5 +$$29.2\pm 1.9$ $55.9\pm 2.2,91.4\pm 0.8$ $59.2\pm 1.1,86.4\pm 0.8$
Gemma 2 9B Untargeted $85.5\pm 1.9$ $79.8\pm 1.1$ $\pm 0.0,30.5\pm 1.7 0.0±0.0$$\pm 0.0,32.6\pm 1.2$
Targeted $68.8\pm 3.8 +$$31.9\pm 2.2$ $63.0\pm 1.9,82.3\pm 1.4$ $61.3\pm 1.2,84.8\pm 0.6$
Table 2: Attack results for all eight settings for Llama-3-8B and Gemma-2-9B on our generated prompt dataset. For individual level, we include results for both activation (left) and deactivation (right) tasks. The mean values and standard deviations are computed based on three independent runs.
Figure 1: Examples of GCG and evaluation losses across iterations for targeted (left) and untargeted (right) population-level suffix attacks.
Example: Targeted Population-level Suffix Attack $x_{1}$: His research surveys Islamic architecture from Morocco to India, revealing regional adaptations and shared aesthetic principles $x_{2}$: Autonomous vehicles rely on sensor fusion, machine learning, and path planning to navigate complex traffic environments $x^{\prime}_{1}$: His research surveys Islamic architecture from Morocco to India, revealing regional adaptations and shared aesthetic principles UAV AMC algorithms
Example: Untargeted Population-level Suffix Attack $x_{1}$: Philosophical skepticism challenges empirical certainty, questioning whether knowledge is possible in the face of doubt $x^{\prime}_{1}$: Philosophical skepticism challenges empirical certainty, questioning whether knowledge is possible in the face of doubt setImageBitmap
Figure 2: Instances of adversarial attacks correspond to the loss curves in Figure 1. Adversarial suffixes are highlighted in red.
4.2 Results Interpretation
We include the complete results for our generated Art & Science dataset in Table 2, and provide the loss curves and adversarial inputs for two representative attack examples in Figure 1 and 2. For tasks on population level, we report the change in overlap ratio (as a percentage) between two sets of $k$ most activated latents, while for those on individual level we report the attack success rate (ASR). The experiment details and results for the AG News dataset are included in Appendix C. In general, our attacks are effective in most cases, and we discuss several important insights below.
Failure Cases
The adversarial attack is only ineffective in the untargeted tasks at the individual level: the SAE latents with lowest activation values cannot be activated by input perturbations. This is most likely due to dead latents (i.e. latents cannot be activated under any circumstances) prevalent in pretrained SAEs, which is aligned with previous findings by Gao et al. (2024).
Untargeted vs. Targeted
Empirically, we find that untargeted attacks are more successful than targeted ones at the population level, primarily because suppressing dominant features is generally easier than constructing specific new ones. However, untargeted attacks are less effective at the individual level, likely because, in targeted settings, the selected neurons are guaranteed to be manipulatable - their exhibited activation status for $x_{2}$ are opposite to $x_{1}$. In contrast, untargeted settings may include a substantial number of SAE latents that are either dead (i.e., never activate) or almost always active. In other words, our results show that as long as an SAE latent represents a meaningful concept (i.e. neither permanently active or inactive), there’s a high chance it could be adversarially manipulated.
Suffix vs. Replacement
At the population level, suffix mode is more effective at controlling the overall SAE activation pattern than replacement mode. This trend is probably due to (1) more recent tokens are likely to have greater impact on both the residual stream and the SAE activations, and (2) appending a short suffix provides more fine-grained and predictable optimization steps on the latent representation space. When perturbing individual SAE features, the effectiveness of two perturbation modes varies across semantic and activation goals, with no significant implications.
4.3 Additional Analyses
Decoupling SAE and LLM Robustness
A potential critique of our evaluation strategy is that it does not fully disentangle the robustness of the SAE from that of the underlying LLM. To address this, we select the number of manipulatable tokens (e.g. suffix length) by examining the LLM’s generations before and after adversarial attacks. Since our user prompts end with an instruction to summarize the preceding sentence (see Section 4.1), the generated outputs reflect the LLM’s self-summarization of the semantic content. Then we can leverage an LLM judge (GPT-4.1-mini) to determine whether the semantic content changes notably after the attacks. We evaluate 100 successful attacks for each unique level of input manipulation, and when the total number of attacks exceed 100, we sample randomly while preserving the balance between different settings and token indices (for replacement mode). For each attack, we take the majority vote among 5 generations as the final decision. Based on the results for the suffix lengths and single-token replacement used in our experiments reported in Table 3, we show that our adversarial attacks are mostly targeting the SAEs, without notably affecting the base LLM. The prompt template and self-summarization examples are included in Appendix E.
Model Dataset Suffix (1 token) Suffix (3 tokens) Replacement (1 token)
Llama 3 8B Art & Science $0.98$ $0.93$ $1.0$
AG News $0.95$ - $0.96$
Gemma 2 9B Art & Science $1.0$ $0.94$ $1.0$
AG News $0.94$ - $0.95$
Table 3: Fraction of adversarial attacks that result in no significant differences in model’s self-summarization, using LLM-as-a-judge (GPT-4.1-mini). During generation, the temperature is set to 0.7 and top_p is set to 0.9.
Robustness Change Across Model Depth
The main results presented in this section focus on a single mid-to-late layer for each model. To assess the generalizability of our findings across different model depths, we apply our population-level attacks under suffix mode to additional layers using open-source pretrained SAE weights. Empirically, we find that our attacks remain effective across layers, despite slight decreases in performance upper bound. The experiment results and interpretations are included in Appendix F.
Transferability of Attacks
In reality, the adversary might want to craft a single perturbed sequence that could be used to attack different models, so it’s important to investigate the transferability of our attacks across different LLMs. Since SAE latents of different models encode different semantic concepts, we only investigate transferability at the population level. As presented in Appendix G, we observe that the difference between average neuron overlap changes achieved by the original attacks and transferred attacks are at most $25.3\%$ for the targeted settings and $43.8\%$ for the untargeted settings, suggesting that although there’s a notable performance gap due to the model transfer, the attacks are still effective.
Deactivating Highly Interpretable SAE Latents from Neuronpedia
In the above experiments, the set of manipulated SAE latents is empirically determined by our datasets (i.e., the most or least activated ones by our inputs). In fact, such evaluation procedure could be reversed: we may instead begin by selecting specific target SAE latents and then assess their robustness by identifying input sequences that strongly activate them. To explore this, we select several SAE latents associated with consistent and human-interpretable semantic concepts from Neuronpedia, and apply the untargeted individual-level attack to top-activating sequences drawn from an external text corpus. Through experiments, we find that our attack could successfully deactivate these highly meaningful SAE latents, with illustrative examples provided in Appendix H.
5.  Discussion and Limitations
The effectiveness of our attacks is fundamentally bounded by compute constraints. All experiments and hyperparameter choices were based on a single 80GB A100 GPU. With access to more GPU memory or increased GCG iterations, even stronger attack performance is likely achievable.
As discussed in section 4.1, we target medium-sized LLMs because smaller models, along with their SAEs, lack the capacity to distinguish semantically unrelated inputs. However, this choice results in slower attacks: for instance, a 50-iteration targeted population-level suffix attack on a longer prompt from the Art & Science dataset takes approximately 10 minutes. In practical applications, users may only need to evaluate the most relevant scenario among our eight settings. Additionally, future work could explore optimizing the attack pipeline to better balance effectiveness and efficiency.
More broadly, we consider our evaluation framework and adversarial attacks as a general methodology for assessing concept-extraction tools for LLMs. While we focus on standard SAEs in this work, the same vulnerabilities likely extend to other variants such as transcoders (Dunefsky et al., 2024) and crosscoders (Lindsey et al., 2024), which similarly lack structural constraints or robustness-aware objectives during training. We leave the exploration of such extensions to future work.
6.  Conclusion
In this work, we investigate the robustness of SAEs under input-level adversarial perturbations and introduce a comprehensive evaluation framework spanning semantic and activation-level objectives. Our experiments show that SAE interpretations are highly vulnerable to minimal input changes, even when the underlying LLM remains semantically stable, raising concerns about their reliability in realistic settings. To advance trustworthy interpretability, we hope our work motivates the development of more robust tools for understanding LLMs, as stability under real-world conditions is essential for aligning model behavior with human expectations.
Appendix A Pseudocode for Adversarial Input-Level Attack
Algorithm 1 Generalized Input-level Attack for SAE
Input: Input token sequence $(x_{1})_{1:l}$, reference input $x_{\rm{ref}}$ (either $x_{1}$ or $x_{2}$), target LLM with the mapping $f_{LLM}:\mathcal{X}\rightarrow\mathcal{H}$, SAE encoding weights $W_{enc}$ and $b_{enc}$, set of modifiable indices $\mathcal{I}$, number of iterations $T$, GCG loss $\mathcal{L}_{gcg}$, evaluation metric $\mathcal{L}_{eval}$, $m$, batch size $B$
$x_{1}^{\prime}\leftarrow\begin{cases}x_{1}&\text{if }\mathcal{I}\subseteq\{1,\dots,l\}\\ \text{Concat}((x_{1})_{1:l},\text{LLM}(x_{1})_{\mathcal{I}})&\text{otherwise}\end{cases}$ $\triangleright$ Initialize $x_{1}^{\prime}$ based on attack mode
for $t=1,\dots,T$ do
  for $i\in\mathcal{I}$ do
   $S_{i}\leftarrow\text{Top-}m(-\nabla_{(x_{1}^{\prime})_{i}}\mathcal{L}_{gcg}(x_{1}^{\prime},x_{\rm{ref}}))$ $\triangleright$ Compute top-$k$ promising token substitutions
  end for
  for $b=1,\dots,B$ do
   $x_{1}^{\prime(b)}\leftarrow x_{1}^{\prime}$ $\triangleright$ Initialize each element within batch
   $x_{1,i}^{\prime(b)}\leftarrow\text{Uniform}(S_{i})$, where $i=\text{Uniform}(\mathcal{I})$ $\triangleright$ Randomly select the token to be replaced
  end for
  $x_{1}^{\prime}\leftarrow\begin{cases}x_{1}^{\prime(b^{*})}$, where $b^{*}=\arg\min_{b}\mathcal{L}_{eval}(x_{1}^{\prime(b)})&\text{if }\mathcal{L}_{eval}(x_{1}^{\prime(b^{*})})<\mathcal{L}_{eval}(x_{1}^{\prime})\\ x_{1}^{\prime}&\text{otherwise}\end{cases}$
end for
Output: Optimized input $x_{1}^{\prime}$
Appendix B Hyperparameters for Different Attacks
Semantic Goal Population Level Individual Level
Suffix Replacement Suffix Replacement
Targeted $T$ $50$ $30$ $10$ $10$
$m$ $300$ $300$ $300$ $300$
$B$ $600$ $200$ $100$ $100$
Untargeted $T$ $10$ $10$ $10$ $10$
$m$ $300$ $300$ $300$ $300$
$B$ $200$ $200$ $100$ $100$
Table B1: Recommended hyperparameters for different types of attacks when running on a 80GB A100 GPU, including the number of iterations $T$, the number of promising tokens considered at each token index $m$, and the batch size $B$. The suffix length is set to 33 under the Targeted-Population setting for the generated Art & Science dataset, and to $1$ in all other cases.
Appendix C Additional Results on AG News Dataset
The complete evaluation results for the AG News dataset are summarized in Table C1. Since the news titles are generally shorter than the prompts in our generated Art & Science dataset, we restrict the suffix length to one token in all settings.
Model Semantic Goal Population Level Individual Level
Neuron Overlap (%) Attack Success Rate (%)
Suffix Replacement Suffix Replacement
Llama 3 8B Untargeted $82.4\pm 1.2$ $77.8\pm 0.7$ $\pm 0.0,24.0\pm 1.0 0.0±0.0$$\pm 0.0,27.9\pm 0.3$
Targeted $64.5\pm 1.8 +$$26.3\pm 0.8$ $60.4\pm 1.3,95.4\pm 0.1$ $57.3\pm 0.7,88.1\pm 0.2$
Gemma 2 9B Untargeted $85.5\pm 0.6$ $73.6\pm 0.6$ $\pm 0.0,28.1\pm 1.4 0.0±0.0$$\pm 0.0,29.3\pm 0.5$
Targeted $56.2\pm 2.0 +$$24.8\pm 1.1$ $62.7\pm 0.9,83.8\pm 0.4$ $64.0\pm 0.5,92.2\pm 0.1$
Table C1: Attack results for all eight settings for Llama-3-8B and Gemma-2-9B on the AG News (300 samples) dataset. Both suffix and replacement modes are restricted to one token. For individual level, we include results for both activation (left) and deactivation (right) tasks. The mean values and standard deviations are computed based on three independent runs.
Appendix D More Examples of Effective Attacks
We provide more examples of effective attacks under different evaluation scenarios in Figure D1, D2, D3, and D4. Samples from both datasets are included.
Figure D1: Examples of population-level attacks. Adversarial tokens are highlighted. We show one example from each dataset for every setting.
Figure D2: Examples of targeted individual-level suffix attacks, including both activation and deactivation tasks.
Figure D3: Examples of targeted individual-level replacement attacks, including both activation and deactivation tasks.
Figure D4: Examples of untargeted individual-level attacks. Empirically, only deactivation attacks can be successful, possibly due to dead latents (explained in section 4.2).
Appendix E Prompt Template for LLM Self-Summarization Comparisons
In Figure E1 and E2, we provide the prompt template along with two examples of model-generated self-summarizations that capture the semantic content of the inputs. The few-shot learning prompt template for using GPT-4.1-mini as the LLM judge is included in Figure E3. The motivations are discussed in section 4.3. To further validate the faithfulness of the LLM judge, we manually label 100 randomly sampled pairs of model generations as ground truth and compare them to the judge’s predictions. The agreement rate reaches $98\%$, with all disagreements being false negatives.
Example 1: Generated Summarization of Original Input User: She paints surreal dreamscapes that draw on ancestral folklore, mythological figures, and personal symbolism from her cultural background. The previous sentence is about Assistant (Llama 3 8B): the artist, and the following is about the art. The artist is a woman, and the art is a painting. The painting is a dream…
Example 1: Generated Summarization of Perturbed Input User: She paints surreal dreamscapes that draw on ancestral folklore, mythological figures, and personal symbolism from her cultural background RadiationLTE regulatory. The previous sentence is about Assistant (Llama 3 8B): the artist’s work, and the following text is about the artist’s life. The artist was born in 1989 in the city of Tiju…
Figure E1: Self-summarization Example 1
Example 2: Generated Summarization of Original Input User: The commemorative statue honors a once-forgotten civil rights activist, reclaiming public space for historical memory. The previous sentence is about Assistant (Gemma 2 9B): the statue of the civil rights activist, the statue is located in the city of Montgomery, Alabama, the statue is a tribute to the civil rights activist…
Example 2: Generated Summarization of Perturbed Input User: The commemorative statue honors a once-forgotten civil rights activist, reclaiming public space for historical memorydifferential geometryDegrees. The previous sentence is about Assistant (Gemma 2 9B): the differential geometry of curves and surfaces. The differential geometry of curves and surfaces is a branch of mathematics that studies the geometry of curves and surfaces. The…
Figure E2: Self-summarization Example 2
Example 1: LLM-as-a-Judge User: You will be given two (potentially truncated) summarizations of a particular sentence, and your goal is to judge whether the two summarizations are focusing on the same main topic. Please answer Yes or No.
1. The previous sentence is about the importance of oral histories in preserving the lived experiences of marginalized groups across generations. Oral histories offer a unique perspective on the past, providing a window into…
2. The previous sentence is about the importance of oral histories in preserving the lived experiences of marginalized groups across generations. Oral histories offer a unique opportunity to hear the voices of those who have…
User: Yes.
User: You will be given two (potentially truncated) summarizations of a particular sentence, and your goal is to judge whether the two summarizations are focusing on the same main topic. Please answer Yes or No.
1. The previous sentence is about the influence of ancient Greek tragedies on contemporary theater. The author argues that the archetypal themes, dramatic irony, and complex character development of ancient…
2. The previous sentence is about the blockchain, which is a distributed ledger technology that allows for secure, transparent, and tamper-proof transactions. It is a decentralized system that does not…
User: No.
User: You will be given two (potentially truncated) summarizations of a particular sentence, and your goal is to judge whether the two summarizations are focusing on the same main topic. Please answer Yes or No.
1. The previous sentence is about the artist, and the following is about the art. The artist is a woman, and the art is a painting. The painting is a dream…
2. The previous sentence is about the artist’s work, and the following text is about the artist’s life. The artist was born in 1989 in the city of Tiju…
Assistant (GPT-4.1-mini): Yes.
Example 2: LLM-as-a-Judge User: You will be given two (potentially truncated) summarizations of a particular sentence, and your goal is to judge whether the two summarizations are focusing on the same main topic. Please answer Yes or No.
1. The previous sentence is about the importance of oral histories in preserving the lived experiences of marginalized groups across generations. Oral histories offer a unique perspective on the past, providing a window into…
2. The previous sentence is about the importance of oral histories in preserving the lived experiences of marginalized groups across generations. Oral histories offer a unique opportunity to hear the voices of those who have…
User: Yes.
User: You will be given two (potentially truncated) summarizations of a particular sentence, and your goal is to judge whether the two summarizations are focusing on the same main topic. Please answer Yes or No.
1. The previous sentence is about the influence of ancient Greek tragedies on contemporary theater. The author argues that the archetypal themes, dramatic irony, and complex character development of ancient…
2. The previous sentence is about the blockchain, which is a distributed ledger technology that allows for secure, transparent, and tamper-proof transactions. It is a decentralized system that does not…
User: No.
User: You will be given two (potentially truncated) summarizations of a particular sentence, and your goal is to judge whether the two summarizations are focusing on the same main topic. Please answer Yes or No.
1. The previous sentence is about the statue of the civil rights activist, the statue is located in the city of Montgomery, Alabama, the statue is a tribute to the civil rights activist…
2. The previous sentence is about the differential geometry of curves and surfaces. The differential geometry of curves and surfaces is a branch of mathematics that studies the geometry of curves and surfaces. The…
Assistant (GPT-4.1-mini): No.
Figure E3: Few-shot learning template for LLM-as-a-judge, corresponding to the previous summarization examples.
Appendix F SAE Robustness Across Model Depth
As discussed in section 4.3, we evaluate our population-level suffix attacks across multiple layers of LLaMA-3-8B (32 layers) and Gemma-2-9B (42 layers) to assess the generalizability of our findings. Table F1, F2 and Figure F1 report the average neuron overlap ratio before and after the attacks on all Art & Science inputs. In Table F2, the Before values are omitted because the initial overlap is always $100\%$ in the untargeted setting.
For targeted suffix attacks, unlike the results presented in Table 2, which quantify overall effectiveness by relative change as a percentage, we report both Before and After values to avoid misleading interpretations: although the relative increase in overlap tends to grow significantly with layer depth, this is largely due to a lower initial overlap. At the same time, the absolute overlap achieved after the attack also decreases with depth, but still remains substantial as reaching approximately $50\%$ in deeper layers. For the untargeted setting, however, an opposite trend is observed as the attacks become more effective when depth increases.
One plausible explanation for these trends is that deeper layers in large language models tend to encode more abstract, task-specific, and distributed representations, making it harder for a short suffix to consistently steer the model toward activating a fixed set of SAE latents. In contrast, earlier and middle layers often retain more localized and compositional features that are easier to manipulate toward a specific goal. Together, these factors contribute to the observed decrease in both initial and post-attack neuron overlap ratios in deeper layers we see in the targeted task. On the other hand, since minor input perturbations can lead to disproportionate changes in model activations, it becomes easier for untargeted attacks to disrupt existing semantic features without the need for precise control.
Consequently, although the upper bound of attack effectiveness in the targeted setting slightly decreases with model depth, this trend is likely driven by representational shifts across LLM layers rather than properties of the SAEs themselves, and we can conclude that our adversarial attacks are generalizable across different model depths.
Model Layer 10 Layer 20 Layer 30 Layer 40
Llama 3 8B Before $41.7$ $32.4$ $21.4$ -
After $57.9$ $56.5$ $51.8$ -
Gemma 2 9B Before $47.2$ $36.1$ $30.1$ $15.7$
After $64.8$ $57.9$ $50.8$ $39.1$
Table F1: Neuron overlaps before and after targeted population-level suffix attacks for different LLM layers. While the upper-bound of attack performance decreases with model depth, the attacks are still effective.
Model Layer 10 Layer 20 Layer 30 Layer 40
Llama 3 8B 81.7-81.7 89.1-89.1 93.2-93.2 -
Gemma 2 9B 78.3-78.3 83.0-83.0 85.5-85.5 93.7-93.7
Table F2: Neuron overlaps before and after untargeted population-level suffix attacks for different LLM layers. Different from the targeted setting, the untargeted attacks are more effective for deeper layers.
(a) Targeted
(b) Untargeted
Figure F1: Population-level suffix attack performance across different layer depths for Llama-3-8B and Gemma-2-9B.
Appendix G Evaluating Attack Transferability
As motivated in section 4.3, we evaluate the transferability of our attacks by directly applying adversarial inputs generated from one model to another. Since the semantic concepts captured by individual SAE latents are not aligned across models, we restrict our analysis to population-level attacks. In Table G1, we report the relative changes in neuron overlap (as percentage values, consistent with Table 2 and C1), along with the absolute differences compared to the original attacks. The results show that, while there is a noticeable performance drop when transferring attacks across models, the transferred adversarial inputs still induce meaningful changes in activation patterns, indicating that the attacks retain a substantial degree of effectiveness.
Model Transfer Semantic Goal Population Level
Neuron Overlap (%)
Suffix Replacement
[…]
Targeted $49.2\quad(\downarrow 25.3)$ $15.8\quad(\downarrow 13.4)$
Gemma 2 9B $\rightarrow$ Llama 3 8B Untargeted $44.1\quad(\downarrow 41.4)$ $40.9\quad(\downarrow 38.9)$
Targeted $46.1\quad(\downarrow 22.7)$ $18.9\quad(\downarrow 13.0)$
Table G1: Performance of transferred attacks at population level. Here we also report the differences in magnitudes of neuron overlap changes between original attacks and transferred attacks. We use the Art & Science dataset for this experiment.
Appendix H Case Study: Manipulating SAE Latents from Neuronpedia
In Figure H1, we present successful adversarial untargeted individual-level replacement attacks that deactivate two SAE latents associated with human-interpretable concepts. These attacks are conducted on specific layers and pretrained SAEs identified by Neuronpedia, which differ from those used in our main experiments. We use top-activating sentences from an external text corpus and leverage Neuronpedia’s token-level activation records to directly extract the corresponding hidden states, eliminating the need for an additional summarization prompt. Empirically, we find that the replacement attack mode is more effective than the suffix mode in this setting. We hypothesize that the high effectiveness of these attacks stems from the dynamic activation behavior of human-interpretable SAE latents, which tend to respond selectively to specific semantic features rather than exhibiting consistently high or low activations. This selectivity makes them more vulnerable to targeted input manipulations.
Neuronpedia Example 1 Layer 29, Neuron #73147 Activated by hurled, fled, fell, … Top-1 Activated Sequence: Washington was stabbed several times but managed to grab a radio which he hurled against a post Deactivating radio: Washington intelligence stabbed several times but managed to grab a radio which he hurled against a post Deactivating hurled: Washington was stabbed Mesozoic times but managed to grab a radio which he hurled against a post Deactivating post: Washington was stabbed several times but managed to grab a radio which he hurled After a post
Neuronpedia Example 2 Layer 35, Neuron #66255 Activated by was, has, is, … Top-1 Activated Sequence: The effect of lard and sunflower oil making part of a cirrhogenic ration with a high content of fat and deficient protein and choline on the level of total and esterified cholesterol and phospholipids in the blood serum and liver was studied Deactivating lipids: The effect of lard and sunflower oil making part of a Swedishrhogenic ration with a high content of fat and deficient protein and choline on the level of total and esterified cholesterol and phospholipids in the blood serum and liver was studied Deactivating blood: The effect of lard and sunflower oil making part of a cirrhogenic ration with a high content of fat and deficient protein and cholineElaboración the level of total and esterified cholesterol and phospholipids in the blood serum and liver was studied Deactivating was: The effect of lard and sunflower oil making part of a cirrhogenic ration with a high content of fat and deficient protein and choline on the level of total and esterified cholesterol and phospholipids inrawDesc blood serum and liver was studied
Figure H1: Examples of successful adversarial attacks that deactivate two highly interpretable SAE latents in their corresponding top-activating sentences. Tokens highlighted in green indicate the specific LLM hidden states passed to the SAE, which Neuronpedia identifies as positions of high activation for the target latent. We show effective adversarial inputs generated via untargeted replacement attacks that suppress activation at these positions, with adversarial tokens highlighted in red.